OpenStudy (anonymous):

What are the possible rational solutions of: f(x) = x^3 - 10x^2 + 16x

OpenStudy (anonymous):

Any help appreciated

OpenStudy (campbell_st):

try x = -2 as you need to factor it 1st \[f(x) = x(x^2 - 10x + 16)\] the others may be obvious from the factored function above.

OpenStudy (anonymous):

So is it then, x(x-2)(x-8)?

OpenStudy (campbell_st):

yep... that seems to make sense... so you should be able to find the rational solutions... and I made a slight error... its not x = -2.... its x = 2

OpenStudy (anonymous):

So, are the solutions -2 and -8 then? Or am I just the slowest person alive? :P

OpenStudy (campbell_st):

to avoid confusion... you need to solve x - 2 = 0 x - 8 = 0 and x = 0 for the solutions...

OpenStudy (anonymous):

so the solutions are +2,+8, and 0. Thank you so much then! :)

OpenStudy (campbell_st):

thats correct...